Use the graph input tool tohelp you answer the following questions. Enter an amount into the Price field tose the quanbity demanded and quanbity supplied at that price. You will otbegraded on any changes you make tothis graph. Graph Input Tool Market for Teapots Price Dowars par taapat) 16 310usntby Supplied100 Demand 24 16 0 50 100 150 200 250 300 350 400 450 500 QUANTITY (Teapots) The equilibrium price in this markt is S per teapot, and the equilibrium quantity is teapots bought and sold per month. Compiete the following tabie by indicating at esch price whether there isa shortage or surplus in the market, the amount of that shortage or surplus, nd whether 바is places upward or downward pressure on prices. Price (Dollars per teapot) 48 32 Shortage or Surplus Amount (Teapots) Shortage or Surplus Pressure
Answer #1

Equilibrium price in the market = $ 40

Equilibrium quantity = 250 teapots

Price Shortage/Surplus Amount Pressure
48 Surplus 70 Downward
32 Shortage 70 Upward

Surplus = Supply - Demand = 300 - 230 = 70

Shortage = Demand - Supply = 270 - 200 = 70